# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ 11.3 The family of examples
Notice that for every $i\in\{1,2\}$ and every $n\geq 1$, the group $G_{\Gamma_{n,i}}$ is isomorphic to the kernel of the homomorphism $\varphi_{n,i}:G_{\Gamma_{i}}\to\mathbb{Z}/n\mathbb{Z}$ that sends $v_{i}$ to $1$ and all other generators to 00. This gives an injective homomorphism $q_{n,i}:G_{\Gamma_{n,i}}\to G_{\Gamma_{i}}$ with finite index image. More precisely, let $\Gamma_{n,i}[1],\dots,\Gamma_{n,i}[n]$ be the $n$ copies of $\Gamma_{i}$ in $\Gamma_{n,i}$, where $\Gamma_{n,i}[j]$ is the copy that contains $w_{n,i}[j]$. Then (up to reordering the $\Gamma_{n,i}[j]$, which we do once and for all), for any vertex $u\in\Gamma_{n,i}[j]$ with $u\neq v_{n,i}$, we have $q_{n,i}(u)=v_{i}^{j-1}\bar{u}v_{i}^{-j+1}$, where $\bar{u}$ denotes the vertex of $\Gamma_{i}$ that corresponds to $u$ (when writing the above equality, we identify vertices of $\Gamma_{i}$ and $\Gamma_{n,i}$ with the corresponding elements of $G_{\Gamma_{i}}$ and $G_{\Gamma_{n,i}}$). And $q_{n,i}(v_{n,i})=v_{i}^{n}$. In particular $q_{n,i}(w_{n,i}[j])=v_{i}^{j-1}w_{i}v_{i}^{-j+1}$. For every $i\in\{1,2\}$ and every $n\geq 1$, the $q_{n,i}$-image of any standard line of $G_{\Gamma_{n,i}}$ is Hausdorff close to a standard line of $G_{\Gamma_{i}}$. In addition $q_{n,i}$ sends parallel standard lines to parallel standard lines, up to finite Hausdorff distance. Thus $q_{n,i}$ induces a map from the vertex set of the extension graph $\Gamma_{n,i}^{e}$ to the vertex set of $\Gamma_{i}^{e}$, which extends to a graph isomorphism $(q_{n,i})_{*}:\Gamma_{n,i}^{e}\to\Gamma_{i}^{e}$. Likewise we have a map $q_{n}:G_{\Lambda_{n}}\to G_{\Lambda}$.
@@ 11.4 Auxiliary facts about star projections
Given a finite simplicial graph $\Gamma$ with $|\operatorname{Out}(G_{\Gamma})|<+\infty$, we collect several facts about certain projections on $G_{\Gamma}$ for later use. Recall that we identify $G_{\Gamma}$ as the 00-skeleton of $X_{\Gamma}$. Given $\mathsf{v}\in V\Gamma^{e}$, take a $\mathsf{v}$-line $\ell$. Then $\ell$ is the 0-skeleton of a convex subcomplex of $X_{\Gamma}$. Then there is nearest point projection $\pi_{\ell}:G_{\Gamma}\to\ell$ sending each point to the nearest point in $\ell$ with respect to the word distance [HW08, Lemma 13.8]. It is known that if $\ell_{1},\ell_{2}$ are standard lines with $\Delta(\ell_{1})=\Delta(\ell_{2})\notin\mathrm{st}(\mathsf{v})$, then $\pi_{\ell}(\ell_{i})$ is a single point for every $i\in\{1,2\}$, and $\pi_{\ell}(\ell_{1})=\pi_{\ell}(\ell_{2})$ (see [Hua17, Lemma 6.2]). This gives a well-defined map $\pi_{\mathsf{v}}:V(\Gamma^{e}\setminus\mathrm{st}(\mathsf{v}))\to\ell$.
@@
Now we prove the fact. Let $x\in G_{\Gamma}$ and $h\in H_{\mathsf{v}}$. Let $y=\pi_{P_{\mathsf{v}}}(x)$. Let $\omega$ be a shortest edge path in the 1-skeleton of $X_{\Gamma}$ connecting the vertices $x$ and $y\in P_{\mathsf{v}}$. Let $x_{0},\dots,x_{n}$ be vertices in $\omega$ such that for $10\leq i\leq n-$ 1, $[x_{i},x_{i+1}]$ is a maximal sub-segment of $\omega$ that is contained in a standard line ($x_{0}=x$ and $x_{n}=y$). Denote the corresponding standard line by $\ell_{i}$. As $y$ can be alternatively characterized as the unique point in $P_{\mathsf{v}}$ such that any hyperplane separating $x$ and $y$ does not cross $C_{\mathsf{v}}$ (see [HW08, Lemmas 13.1 and 13.8]), $y=\pi_{P_{\mathsf{v}}}(x)$ if and only if for any $\omega$ and $(\ell_{i})$ as above, $\Delta(\ell_{i})\notin\mathrm{st}(\mathsf{v})$ for any $10\leq i\leq n-$ 1.
For $10\leq i\leq n-$ 1, let $\ell^{\prime}_{i}$ be the standard line with $h(\ell_{i})=\ell^{\prime}_{i}$ and let $\omega^{\prime}_{i}$ be a geodesic segment in $\ell^{\prime}_{i}$ from $h(x_{i})$ to $h(x_{i+1})$. Let $\omega^{\prime}$ be the concatenation of all the $\omega^{\prime}_{i}$ with $10\leq i\leq n-$ 1. By considering the automorphism $h_{*}:\Gamma^{e}\to\Gamma^{e}$ induced by the flat-preserving bijection $h$, we see that $\Delta(\ell^{\prime}_{i})\notin\mathrm{st}(\mathsf{v})$ for any $10\leq i\leq n-$ 1. Thus none of the hyperplanes that have nonempty intersection with $\omega^{\prime}$ will intersect $C_{\mathsf{v}}$. As any hyperplane separating $h(x)$ and $h(y)$ is dual to an edge in $\omega^{\prime}$, this hyperplane does not cross $C_{\mathsf{v}}$. Thus $h(y)=\pi_{P_{\mathsf{v}}}(h(x))$ by the previous paragraph. ∎
@@ 11.5 Conclusion
Let $G$ be a non-cyclic right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$. Let $\rho$ be a quasi-action of $H$ on $G$, and let $\alpha:H\curvearrowright G$ be the unique $H$-action on $G$ by flat-preserving bijections which is equivalent to $\rho$ (see Corollary 10.3). Then there exists $C\geq 0$ (depending on $\rho$) such that for every $\mathsf{v}\in V\Gamma^{e}$ and every subgroup $H^{\prime}_{\mathsf{v}}\subseteq H_{\mathsf{v}}$, if some $H^{\prime}_{\mathsf{v}}$-orbit of the factor action $\alpha_{\mathsf{v}}$ is finite, then every $H^{\prime}_{\mathsf{v}}$-orbit for this action has cardinality at most $C$.
@@
Fix $n_{0}\in\mathcal{C}$. Since $U^{\prime}_{n_{0}}$ is quasi-isometric to $G_{\Lambda}$, we can apply Proposition 10.4 and get a proper, cobounded action $\alpha_{1}$ of $\mathfrak{G}$ on $G_{\Lambda}$ by flat-preserving bijective quasi-isometries with uniform constants. This gives, for every $n\in\mathcal{C}$, an action by flat-preserving bijections $\alpha_{1,n}:U^{\prime}_{n}\curvearrowright G_{\Lambda}$ through the lattice embedding $U^{\prime}_{n}\to\mathfrak{G}$. We apply Proposition 11.7 to the action $\alpha_{1}:\mathfrak{G}\curvearrowright G_{\Lambda}$, and let $C$ be the resulting constant. On the other hand, for every $n\in\mathcal{C}$, the group $U^{\prime}_{n}$ has a properly discontinuous cocompact action on $X_{2n,1}\times X_{2n,2}$. Let $q_{2n}:X_{2n,1}\times X_{2n,2}\to G_{\Lambda}$ be as in the previous section, after identifying the 00-skeleton of $X_{2n,1}\times X_{2n,2}$ with $G_{\Lambda_{2n}}$. Then $q_{2n}$ gives a quasi-action of $U^{\prime}_{n}$ on $G_{\Lambda}$, which is equivalent to a unique action by flat-preserving bijections $\alpha_{2,n}:U^{\prime}_{n}\curvearrowright G_{\Lambda}$ (see Corollary 10.3). Note that both $\alpha_{1,n}$ and $\alpha_{2,n}$ are proper and cobounded actions of $U^{\prime}_{n}$ on $G_{\Lambda}$ by quasi-isometries with uniform constants. Thus $\alpha_{1,n}$ and $\alpha_{2,n}$ are quasi-conjugate (as any two proper and cobounded quasi-actions of the same finitely generated group on the same space are quasi-conjugate). By Theorem 10.2, $\alpha_{1,n}$ and $\alpha_{2,n}$ are actually conjugate via a flat-preserving bijection of $G_{\Lambda}$.
@@
Let $\{\alpha_{2,n,\mathsf{v}}:U^{\prime}_{n,\mathsf{v}}\curvearrowright Z_{\mathsf{v}}\}_{\mathsf{v}\in V\Lambda^{e}}$ be the collection of factor actions for $\alpha_{2,n}:U^{\prime}_{n}\curvearrowright G_{\Lambda}$. We observe that for every subgroup $U^{\prime\prime}_{n,\mathsf{v}}\subseteq U^{\prime}_{n,\mathsf{v}}$, every finite orbit of $(\alpha_{2,n,\mathsf{v}})_{|U^{\prime\prime}_{n,\mathsf{v}}}$ has cardinality at most $C$: indeed, this follows from the same property for the action $\alpha_{1,n,\mathsf{v}}$, which comes from our choice of constant $C$, and from the fact that the actions $\alpha_{1,n}$ and $\alpha_{2,n}$ are conjugate via a flat-preserving bijection. In the rest of the proof, we will show that as $n$ becomes larger and larger, we can find some $\mathsf{v}\in V\Lambda^{e}$ and some subgroup $U^{\prime\prime}_{n,\mathsf{v}}$ of $U^{\prime}_{n,\mathsf{v}}$ acting on $Z_{\mathsf{v}}$ with finite orbits of larger and larger size, which will be a contradiction. Consider the simple subgroup $V_{n}$ of $U_{n}$ as above. Then $V_{n}\subseteq U^{\prime}_{n}$. We view $T_{n}\times T_{n}$ as a $V_{n}$-invariant subcomplex of $X_{2n,1}\times X_{2n,2}$ via the embedding $\theta_{n}$ provided by Corollary 11.4. A vertical (resp. horizontal) $T_{n}$-copy in $X_{2n,1}\times X_{2n,2}$ is the $\theta_{n}$-image of $\{z\}\times T_{n}$ (resp. $T_{n}\times\{z\}$) for some vertex $z$ in the first (resp. second) tree factor. Let $\ell\subset X_{2n,1}\times X_{2n,2}$ be a standard line of type $v_{2n,1}$ such that $\ell$ intersects $T_{n}\times T_{n}$ in a vertex $x\in X_{2n,1}\times X_{2n,2}$. Let $p_{1}:T_{n}\times T_{n}\to T_{n}$ be the projection to the first factor. Let $V_{n,x}$ be the $V_{n}$-stabilizer of $p_{1}(x)$ with respect to the action of $V_{n}$ on the first factor of $T_{n}\times T_{n}$. We claim that each element in $V_{n,x}$ sends $\ell$ to a standard line which is parallel to $\ell$. Indeed, by Corollary 11.4(5), for any $g\in V_{n,x}$, edges in $g(\ell)$ are also labeled by $v_{2n,1}$. On the other hand, $x$ and $g(x)$ are connected by an edge path in a vertical $T_{n}$-copy whose edge labels commute with $v_{2n,1}$, thus $\ell$ and $g(\ell)$ are parallel. Let $P_{\ell}$ be the parallel set of $\ell$ in $G_{\Lambda_{2n}}$. We write $P_{\ell}=g_{0}G_{\mathrm{st}(v_{2n,1})}$ and assume without loss of generality that $x\in g_{0}G_{\mathrm{lk}(v_{2n,1})}$. By the previous paragraph, $V_{n,x}$ stabilizes both $P_{\ell}$ and $g_{0}G_{\mathrm{lk}(v_{2n,1})}$. Let $T_{n,x}^{h}$ be the horizontal $T_{n}$-copy that contains $x$. Let $\{\ell_{1},\dots,\ell_{k}\}$ be the set of standard lines in $X_{2n,1}\times X_{2n,2}$ passing through $x$ whose intersection with $T_{n,x}^{h}$ contains at least one edge. By Corollary 11.4(4), we have $k=n$, and the types of the lines $\ell_{1},\dots,\ell_{n}$ are contained in $\{w_{2n,1}[1],\ldots,w_{2n,1}[2n]\}$. For every $i\in\{1,\dots,n\}$, let $e_{i}\subseteq\ell_{i}$ be an edge based at $x$ and contained in $\ell_{i}\cap T_{n,x}^{h}$. Since the action of $U_{n}$ on $X_{2n,1}\times X_{2n,2}$ sends standard flat to standard flat (Corollary 11.4), for every $g\in V_{n,x}$ and every $i\in\{1,\dots,n\}$, the image $g\ell_{i}$ is a standard line, and $g(e_{i})$ is an edge based at $gx$ contained in $g\ell_{i}\cap T_{n,gx}^{h}$. It thus follows from Lemma 11.3 and Corollary 11.4(4) that, up to permuting the lines $\ell_{i}$, the set of all types of $g\ell_{1}$ as $g$ varies in $V_{n,x}$, is a subset of $\{w_{2n,1}[1],\dots,w_{2n,1}[2n]\}$ of cardinality at least $n/4$.
@@
Let $\ell^{\prime}_{1}$ (resp. $\ell^{\prime}$) be a standard line in $G_{\Lambda}$ at finite Hausdorff distance from $q_{2n}(\ell_{1})$ (resp. $q_{2n}(\ell)$). Let $\mathsf{v}=\Delta(\ell^{\prime})$. We consider the action $\alpha_{2,n,\mathsf{v}}:U^{\prime}_{n,\mathsf{v}}\curvearrowright Z_{\mathsf{v}}$. The fact that every element of $V_{n,x}$ sends $\ell$ to a parallel line ensures that $V_{n,x}\subset U^{\prime}_{n,\mathsf{v}}$. Notice that for every $g\in V_{n,x}$, we have $(q_{2n})_{\ast}(\Delta(g\ell_{1}))=\Delta(g\ell^{\prime}_{1})$. Thus, using Lemma 11.6, we deduce that the projections $\pi_{\mathsf{v}}((q_{2n})_{*}(\Delta(g\ell_{1})))$, as $g$ varies in $V_{n,x}$, form exactly one orbit of the action $V_{n,x}\curvearrowright Z_{\mathsf{v}}$ under $\alpha_{2,n,\mathsf{v}}$. As the types of the lines $g\ell_{1}$, with $g$ varying in $V_{n,x}$, form a subset of $\{w_{2n,1}[1],\ldots,w_{2n,1}[2n]\}$ of cardinality at least $n/4$, Lemma 11.5(1) implies that the cardinality of the set $\{\pi_{\mathsf{v}}((q_{2n})_{*}(\Delta(g\ell_{1})))\}_{g\in V_{n,x}}$ is at least $n/4$. In addition, since $\Delta(\ell_{1})$ and $\Delta(\ell)$ are not adjacent in $\Lambda_{2n}^{e}$, and $(q_{2n})_{\ast}$ is an isomorphism, we have $\Delta(\ell^{\prime}_{1})\notin\mathrm{st}(\Delta(\ell^{\prime}))$. We can thus apply Lemma 11.5(2) and deduce that the above set (which as we explained is exactly one orbit of $V_{n,x}$ for the action $\alpha_{2,n,\mathsf{v}}$) is finite. We have thus found a subgroup of $U^{\prime}_{n,\mathsf{v}}$ whose action on $Z_{\mathsf{v}}$ via $\alpha_{2,n,\mathsf{v}}$ has a finite orbit of size at least $n/4$, as desired. ∎

### v2

@@ 11.3 The family of examples
Notice that for every $i\in\{1,2\}$ and every $n\geq 1$, the group $G_{\Gamma_{n,i}}$ is isomorphic to the kernel of the homomorphism $\varphi_{n,i}:G_{\Gamma_{i}}\to\mathbb{Z}/n\mathbb{Z}$ that sends $v_{i}$ to $1$ and all other generators to 0. This gives an injective homomorphism $q_{n,i}:G_{\Gamma_{n,i}}\to G_{\Gamma_{i}}$ with finite index image. More precisely, let $\Gamma_{n,i}[1],\dots,\Gamma_{n,i}[n]$ be the $n$ copies of $\Gamma_{i}$ in $\Gamma_{n,i}$, where $\Gamma_{n,i}[j]$ is the copy that contains $w_{n,i}[j]$. Then (up to reordering the $\Gamma_{n,i}[j]$, which we do once and for all), for any vertex $u\in\Gamma_{n,i}[j]$ with $u\neq v_{n,i}$, we have $q_{n,i}(u)=v_{i}^{j-1}\bar{u}v_{i}^{-j+1}$, where $\bar{u}$ denotes the vertex of $\Gamma_{i}$ that corresponds to $u$ (when writing the above equality, we identify vertices of $\Gamma_{i}$ and $\Gamma_{n,i}$ with the corresponding elements of $G_{\Gamma_{i}}$ and $G_{\Gamma_{n,i}}$). And $q_{n,i}(v_{n,i})=v_{i}^{n}$. In particular $q_{n,i}(w_{n,i}[j])=v_{i}^{j-1}w_{i}v_{i}^{-j+1}$.
For every $i\in\{1,2\}$ and every $n\geq 1$, the $q_{n,i}$-image of any standard line of $G_{\Gamma_{n,i}}$ is Hausdorff close to a standard line of $G_{\Gamma_{i}}$. In addition $q_{n,i}$ sends parallel standard lines to parallel standard lines, up to finite Hausdorff distance. Thus $q_{n,i}$ induces a map from the vertex set of the extension graph $\Gamma_{n,i}^{e}$ to the vertex set of $\Gamma_{i}^{e}$, which extends to a graph isomorphism $(q_{n,i})_{*}:\Gamma_{n,i}^{e}\to\Gamma_{i}^{e}$. Likewise we have a map $q_{n}:G_{\Lambda_{n}}\to G_{\Lambda}$.
@@ 11.4 Auxiliary facts about star projections
Given a finite simplicial graph $\Gamma$ with $|\operatorname{Out}(G_{\Gamma})|<+\infty$, we collect several facts about certain projections on $G_{\Gamma}$ for later use. Recall that we identify $G_{\Gamma}$ as the 0-skeleton of $X_{\Gamma}$. Given $\mathsf{v}\in V\Gamma^{e}$, take a $\mathsf{v}$-line $\ell$. Then $\ell$ is the 0-skeleton of a convex subcomplex of $X_{\Gamma}$. Then there is nearest point projection $\pi_{\ell}:G_{\Gamma}\to\ell$ sending each point to the nearest point in $\ell$ with respect to the word distance [HW08, Lemma 13.8]. It is known that if $\ell_{1},\ell_{2}$ are standard lines with $\Delta(\ell_{1})=\Delta(\ell_{2})\notin\mathrm{st}(\mathsf{v})$, then $\pi_{\ell}(\ell_{i})$ is a single point for every $i\in\{1,2\}$, and $\pi_{\ell}(\ell_{1})=\pi_{\ell}(\ell_{2})$ (see [Hua17, Lemma 6.2]). This gives a well-defined map $\pi_{\mathsf{v}}:V(\Gamma^{e}\setminus\mathrm{st}(\mathsf{v}))\to\ell$.
@@
Now we prove the fact. Let $x\in G_{\Gamma}$ and $h\in H_{\mathsf{v}}$. Let $y=\pi_{P_{\mathsf{v}}}(x)$. Let $\omega$ be a shortest edge path in the 1-skeleton of $X_{\Gamma}$ connecting the vertices $x$ and $y\in P_{\mathsf{v}}$. Let $x_{0},\dots,x_{n}$ be vertices in $\omega$ such that for $0\leq i\leq n-1$, $x_{i},x_{i+1}]$ is a maximal sub-segment of $\omega$ that is contained in a standard line ($x_{0}=x$ and $x_{n}=y$). Denote the corresponding standard line by $\ell_{i}$. As $y$ can be alternatively characterized as the unique point in $P_{\mathsf{v}}$ such that any hyperplane separating $x$ and $y$ does not cross $C_{\mathsf{v}}$ (see [HW08, Lemmas 13.1 and 13.8]), $y=\pi_{P_{\mathsf{v}}}(x)$ if and only if for any $\omega$ and $(\ell_{i})$ as above, $\Delta(\ell_{i})\notin\mathrm{st}(\mathsf{v})$ for any $0\leq i\leq n-1$.
For $0\leq i\leq n-1$, let $\ell^{\prime}_{i}$ be the standard line with $h(\ell_{i})=\ell^{\prime}_{i}$ and let $\omega^{\prime}_{i}$ be a geodesic segment in $\ell^{\prime}_{i}$ from $h(x_{i})$ to $h(x_{i+1})$. Let $\omega^{\prime}$ be the concatenation of all the $\omega^{\prime}_{i}$ with $0\leq i\leq n-1$. By considering the automorphism $h_{*}:\Gamma^{e}\to\Gamma^{e}$ induced by the flat-preserving bijection $h$, we see that $\Delta(\ell^{\prime}_{i})\notin\mathrm{st}(\mathsf{v})$ for any $0\leq i\leq n-1$. Thus none of the hyperplanes that have nonempty intersection with $\omega^{\prime}$ will intersect $C_{\mathsf{v}}$. As any hyperplane separating $h(x)$ and $h(y)$ is dual to an edge in $\omega^{\prime}$, this hyperplane does not cross $C_{\mathsf{v}}$. Thus $h(y)=\pi_{P_{\mathsf{v}}}(h(x))$ by the previous paragraph. ∎
@@ 11.5 Conclusion
Let $G$ be a non-cyclic right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$. Let $\rho$ be a quasi-action of $H$ on $G$, and let $\alpha:H\curvearrowright G$ be the unique $H$-action on $G$ by flat-preserving bijections which is equivalent to $\rho$ (see Corollary 10.3).
Then there exists $C\geq 0$ (depending on $\rho$) such that for every $\mathsf{v}\in V\Gamma^{e}$ and every subgroup $H^{\prime}_{\mathsf{v}}\subseteq H_{\mathsf{v}}$, if some $H^{\prime}_{\mathsf{v}}$-orbit of the factor action $\alpha_{\mathsf{v}}$ is finite, then every $H^{\prime}_{\mathsf{v}}$-orbit for this action has cardinality at most $C$.
@@
Fix $n_{0}\in\mathcal{C}$. Since $U^{\prime}_{n_{0}}$ is quasi-isometric to $G_{\Lambda}$, we can apply Proposition 10.4 and get a proper, cobounded action $\alpha_{1}$ of $\mathfrak{G}$ on $G_{\Lambda}$ by flat-preserving bijective quasi-isometries with uniform constants. This gives, for every $n\in\mathcal{C}$, an action by flat-preserving bijections $\alpha_{1,n}:U^{\prime}_{n}\curvearrowright G_{\Lambda}$ through the lattice embedding $U^{\prime}_{n}\to\mathfrak{G}$. We apply Proposition 11.8 to the action $\alpha_{1}:\mathfrak{G}\curvearrowright G_{\Lambda}$, and let $C$ be the resulting constant.
On the other hand, for every $n\in\mathcal{C}$, the group $U^{\prime}_{n}$ has a properly discontinuous cocompact action on $X_{2n,1}\times X_{2n,2}$. Let $q_{2n}:X_{2n,1}\times X_{2n,2}\to G_{\Lambda}$ be as in the previous section, after identifying the 0-skeleton of $X_{2n,1}\times X_{2n,2}$ with $G_{\Lambda_{2n}}$. Then $q_{2n}$ gives a quasi-action of $U^{\prime}_{n}$ on $G_{\Lambda}$, which is equivalent to a unique action by flat-preserving bijections $\alpha_{2,n}:U^{\prime}_{n}\curvearrowright G_{\Lambda}$ (see Corollary 10.3). Note that both $\alpha_{1,n}$ and $\alpha_{2,n}$ are proper and cobounded actions of $U^{\prime}_{n}$ on $G_{\Lambda}$ by quasi-isometries with uniform constants. Thus $\alpha_{1,n}$ and $\alpha_{2,n}$ are quasi-conjugate (as any two proper and cobounded quasi-actions of the same finitely generated group on the same space are quasi-conjugate). By Theorem 10.2, $\alpha_{1,n}$ and $\alpha_{2,n}$ are actually conjugate via a flat-preserving bijection of $G_{\Lambda}$.
@@
Let $\{\alpha_{2,n,\mathsf{v}}:U^{\prime}_{n,\mathsf{v}}\curvearrowright Z_{\mathsf{v}}\}_{\mathsf{v}\in V\Lambda^{e}}$ be the collection of factor actions for $\alpha_{2,n}:U^{\prime}_{n}\curvearrowright G_{\Lambda}$. We observe that for every subgroup $U^{\prime\prime}_{n,\mathsf{v}}\subseteq U^{\prime}_{n,\mathsf{v}}$, every finite orbit of $(\alpha_{2,n,\mathsf{v}})_{|U^{\prime\prime}_{n,\mathsf{v}}}$ has cardinality at most $C$: indeed, this follows from the same property for the action $\alpha_{1,n,\mathsf{v}}$, which comes from our choice of constant $C$, and from the fact that the actions $\alpha_{1,n}$ and $\alpha_{2,n}$ are conjugate via a flat-preserving bijection. In the rest of the proof, we will show that as $n$ becomes larger and larger, we can find some $\mathsf{v}\in V\Lambda^{e}$ and some subgroup $U^{\prime\prime}_{n,\mathsf{v}}$ of $U^{\prime}_{n,\mathsf{v}}$ acting on $Z_{\mathsf{v}}$ with finite orbits of larger and larger size, which will be a contradiction.
Consider the simple subgroup $V_{n}$ of $U_{n}$ as above. Then $V_{n}\subseteq U^{\prime}_{n}$. We view $T_{n}\times T_{n}$ as a $V_{n}$-invariant subcomplex of $X_{2n,1}\times X_{2n,2}$ via the embedding $\theta_{n}$ provided by Corollary 11.5. A vertical (resp. horizontal) $T_{n}$-copy in $X_{2n,1}\times X_{2n,2}$ is the $\theta_{n}$-image of $\{z\}\times T_{n}$ (resp. $T_{n}\times\{z\}$) for some vertex $z$ in the first (resp. second) tree factor. Let $\ell\subset X_{2n,1}\times X_{2n,2}$ be a standard line of type $v_{2n,1}$ such that $\ell$ intersects $T_{n}\times T_{n}$ in a vertex $x\in X_{2n,1}\times X_{2n,2}$. Let $p_{1}:T_{n}\times T_{n}\to T_{n}$ be the projection to the first factor. Let $V_{n,x}$ be the $V_{n}$-stabilizer of $p_{1}(x)$ with respect to the action of $V_{n}$ on the first factor of $T_{n}\times T_{n}$.
We claim that each element in $V_{n,x}$ sends $\ell$ to a standard line which is parallel to $\ell$. Indeed, by Corollary 11.5(5), for any $g\in V_{n,x}$, edges in $g(\ell)$ are also labeled by $v_{2n,1}$. On the other hand, $x$ and $g(x)$ are connected by an edge path in a vertical $T_{n}$-copy whose edge labels commute with $v_{2n,1}$, thus $\ell$ and $g(\ell)$ are parallel.
Let $P_{\ell}$ be the parallel set of $\ell$ in $G_{\Lambda_{2n}}$. We write $P_{\ell}=g_{0}G_{\mathrm{st}(v_{2n,1})}$ and assume without loss of generality that $x\in g_{0}G_{\mathrm{lk}(v_{2n,1})}$. By the previous paragraph, $V_{n,x}$ stabilizes both $P_{\ell}$ and $g_{0}G_{\mathrm{lk}(v_{2n,1})}$.
Let $T_{n,x}^{h}$ be the horizontal $T_{n}$-copy that contains $x$. Let $\{\ell_{1},\dots,\ell_{k}\}$ be the set of standard lines in $X_{2n,1}\times X_{2n,2}$ passing through $x$ whose intersection with $T_{n,x}^{h}$ contains at least one edge. By Corollary 11.5(4), we have $k=n$, and the types of the lines $\ell_{1},\dots,\ell_{n}$ are contained in $\{w_{2n,1}[1],\ldots,w_{2n,1}[2n]\}$. For every $i\in\{1,\dots,n\}$, let $e_{i}\subseteq\ell_{i}$ be an edge based at $x$ and contained in $\ell_{i}\cap T_{n,x}^{h}$. Since the action of $U_{n}$ on $X_{2n,1}\times X_{2n,2}$ sends standard flat to standard flat (Corollary 11.5), for every $g\in V_{n,x}$ and every $i\in\{1,\dots,n\}$, the image $g\ell_{i}$ is a standard line, and $g(e_{i})$ is an edge based at $gx$ contained in $g\ell_{i}\cap T_{n,gx}^{h}$. It thus follows from Lemma 11.4 and Corollary 11.5(4) that, up to permuting the lines $\ell_{i}$, the set of all types of $g\ell_{1}$ as $g$ varies in $V_{n,x}$, is a subset of $\{w_{2n,1}[1],\dots,w_{2n,1}[2n]\}$ of cardinality at least $n/4$.
Let $\ell^{\prime}_{1}$ (resp. $\ell^{\prime}$) be a standard line in $G_{\Lambda}$ at finite Hausdorff distance from $q_{2n}(\ell_{1})$ (resp. $q_{2n}(\ell)$). Let $\mathsf{v}=\Delta(\ell^{\prime})$. We consider the action $\alpha_{2,n,\mathsf{v}}:U^{\prime}_{n,\mathsf{v}}\curvearrowright Z_{\mathsf{v}}$. The fact that every element of $V_{n,x}$ sends $\ell$ to a parallel line ensures that $V_{n,x}\subset U^{\prime}_{n,\mathsf{v}}$.
@@
Notice that for every $g\in V_{n,x}$, we have $(q_{2n})_{\ast}(\Delta(g\ell_{1}))=\Delta(g\ell^{\prime}_{1})$. Thus, using Lemma 11.7, we deduce that the projections $\pi_{\mathsf{v}}((q_{2n})_{*}(\Delta(g\ell_{1})))$, as $g$ varies in $V_{n,x}$, form exactly one orbit of the action $V_{n,x}\curvearrowright Z_{\mathsf{v}}$ under $\alpha_{2,n,\mathsf{v}}$. As the types of the lines $g\ell_{1}$, with $g$ varying in $V_{n,x}$, form a subset of $\{w_{2n,1}[1],\ldots,w_{2n,1}[2n]\}$ of cardinality at least $n/4$, Lemma 11.6(1) implies that the cardinality of the set $\{\pi_{\mathsf{v}}((q_{2n})_{*}(\Delta(g\ell_{1})))\}_{g\in V_{n,x}}$ is at least $n/4$. In addition, since $\Delta(\ell_{1})$ and $\Delta(\ell)$ are not adjacent in $\Lambda_{2n}^{e}$, and $(q_{2n})_{\ast}$ is an isomorphism, we have $\Delta(\ell^{\prime}_{1})\notin\mathrm{st}(\Delta(\ell^{\prime}))$. We can thus apply Lemma 11.6(2) and deduce that the above set (which as we explained is exactly one orbit of $V_{n,x}$ for the action $\alpha_{2,n,\mathsf{v}}$) is finite. We have thus found a subgroup of $U^{\prime}_{n,\mathsf{v}}$ whose action on $Z_{\mathsf{v}}$ via $\alpha_{2,n,\mathsf{v}}$ has a finite orbit of size at least $n/4$, as desired. ∎